OpenStudy (anonymous):

Would 38% of 8,000 be 3800, leaving the remainder of 4200?

OpenStudy (anonymous):

Here is the actually question, I keep putting 4200 in nd it is saying it is wrong.... 4. At the Boston Children's Museum, it was estimated 38% of all visitors are from in state. On Saturday, 8,000 people attended the museum. What is the number of out of state people in attendance? (Points : 1)

OpenStudy (callisto):

Well... 38% of 8000 is NOT 3800... 38% of 1000 = 380 38% of 8000 = 8 (38% of 1000) = 8 (380) =?

OpenStudy (anonymous):

3040?

OpenStudy (anonymous):

So the answer to the question would be 4960

OpenStudy (callisto):

Yes!~

OpenStudy (anonymous):

So how do I convert the 38% to a whole number again? I read your explanation a few times and still not quite getting it.

OpenStudy (anonymous):

I am one of the few people that actually likes to learn how to get the answer LOL

OpenStudy (callisto):

Oh.. sorry. 38% = 38/100 = 0.38 So, 38% of 8000 = 8000 (0.38)

OpenStudy (mertsj):

If 38% were in state, then 62% were out of state. Multiply .62 times 8000 and that is your answer.